Ibuprofen is an anti-inflammatory with the formula C13H18O2. How many grams of ibuprofen are in 0.495 mol? How many moles of carbon are in 10.0 g of ibuprofen? How many moles of ibuprofen contain 1.15×1023 atoms of carbon?
Answer #1

mass of ibuprofen = mole * molar mass of ibuprofen

molar mass of ibuprofen = 206.3 grams/mole

= 0.495 * 206.3

= 102.1185 grams

moles of ibuprofen in 10 grams = mass / molar mass

= 10.0 / 206.3

= 0.0485

formula of ibuprofen = C13H18O2

1 mole have 13 carbon

0.0485 mole will have = 13 * 0.0485

= 0.6305 moles of carbon

Now for third part

moles of carbon = number of carbon atoms / avagadro number

= 1.15*10^23 / (6.022*10^23)

= 0.191

moles of ibuprofen = 1/13 * moles of C

1/13 * 0.191

= 0.0147 moles

any query please comment

if satisfied please rate it thanks